OpenStudy (anonymous):

Imagine you are teaching a fellow student how to solve: 3 − 5(x + 2) = 3(x + 1) + 6 In your own words, explain the process for solving this equation. Please include your work at each step of the process along with the final answer.

hero (hero):

If I did it, I would scare the student to death first by posting one of my solutions. Then afterward, I would show them the easy way.

OpenStudy (anonymous):

first multiply the brackets 3-5x-10=3x+3+6 -7-5x=3x+9 -16=8x x=-2 if u have any doubt ask